OpenStudy (anonymous):

Which phrase represents the algebraic expression x/3 + 5? five divided by a number, increased by three five greater than a number divided into three the sum of three and a number divided by five five more than the quotient of a number and three

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

five greater than a number divided by 3

OpenStudy (anonymous):

how did you get that answer

OpenStudy (anonymous):

i got the answer thanks

OpenStudy (anonymous):

a number= x /3= divided by 3 +5= five greater :)

OpenStudy (anonymous):

can you answer this one

OpenStudy (anonymous):

Which operation is associated with Reese winning one fewer skateboarding trophy than Vince? addition multiplication subtraction division